how do i turn this equation into standard form
y=1/6+3

Answer by checkley77(9411)

y=1/3+3
y=(1+9)/3
y=10/3 Answer.
Unless you left out the x term.
y=1/3x+3
y=x/3+3 This would be the standard for a line equation.